OpenStudy (anonymous):

A point Q on a segment with endpoints A (2, -1) and C (4, 2) partitions the segment in a 3:1 ratio. Find Q

OpenStudy (anonymous):

k=3/4 x=x1+k(x2-x1) =2+3/4(4-2) =2+3/2 =7/2 y=y1+k(y2-y1) =-1+3/4(2+1) =-1+9/4 =5/4 Therefore, Q is (7/2,5/4).